# fourierlargesymbols147 thanks: Present address for Q. H. Tran: Quantum Laboratory, Fujitsu Research, Fujitsu Limited

### v4

@@ Abstract
The echo state property (ESP) represents a fundamental concept in the reservoir computing (RC) framework that ensures output-only training of reservoir networks by being agnostic to the initial states and far past inputs. However, the traditional definition of ESP does not describe possible non-stationary systems in which statistical properties evolve. To address this issue, we introduce two new categories of ESP: non-stationary ESP, designed for potentially non-stationary systems, and subspace/subset ESP, designed for systems whose subsystems have ESP. Following the definitions, we numerically demonstrate the correspondence between non-stationary ESP in the quantum reservoir computer (QRC) framework with typical Hamiltonian dynamics and input encoding methods using non-linear autoregressive moving-average (NARMA) tasks. We also confirm the correspondence by computing linear/non-linear memory capacities that quantify input-dependent components within reservoir states. Our study presents a new understanding of the practical design of QRC and other possibly non-stationary RC systems in which non-stationary systems and subsystems are exploited.
@@ I Introduction
In this paper, we define a two extensions of traditional ESP. One extension is non-stationary ESP, which requires finite variance output signals relative to initial-state difference decay. Another extension is subset/subspace non-stationary ESP, which focuses on a situation in which a part of the system has a non-stationary ESP, leaving the entire system possibly initial-state dependent. Our analysis includes a numerical study of non-stationary ESP and subset non-stationary ESP in a typical QRC setup with a specific type of Hamiltonian system.
@@
  • Defined non-stationary and subspace/subset versions of ESP, which could be practical for QRC and other non-conventional systems.
@@ II.1 Non-stationary ESP
((a)) Non-stationary ESP
((b)) Traditional ESP with input-invariant fixed point
Figure 2: Schematics of two types of ESPs. State space is illustrated as a circle, and different states are illustrated as red arrows. (a) Non-stationary ESP. State difference decays, but the state space remains finite. (b) With traditional ESP but not with the non-stationary ESP. The state difference decays as fast as the state decay. Schematics show representative cases where two states compared in ESP calculation are on the same shrinking state space for simplicity.
@@
$\mathrm{E}_{w}^{t}(\{\mathbf{u}_{\tau}\})$ $\equiv\frac{1}{w}\sum_{k=0}^{w-1}\mathbf{u}_{t-k}$ (3)
$\mathrm{Var}_{w}^{t}(\{\mathbf{u}_{\tau}\})$ $\equiv{\mathrm{E}}_{w}^{t}\left(\{\|\mathbf{u}_{\tau}-\mathrm{E}_{w}^{t}(\{\mathbf{u}_{\tau}\})\|^{2}\}\right)$
$\overline{\mathrm{E}}_{w}^{t}(f;s_{0})$ $\equiv\frac{1}{w}\sum_{k=0}^{w-1}f(\{\mathbf{u}_{\tau}\}_{\tau\leq t-k};s_{0})$
@@ II.2 Subspace and subset non-stationary ESP
((a)) Subspace ESP
((b)) Subset ESP
Figure 3: (a) In the subspace ESP, the full state space $\mathcal{S}$ has an invariant subspace $\mathcal{S^{\prime}}\subseteq\mathcal{S}$ under the input-driven system dynamics $f$. Furthermore, $\mathcal{S^{\prime}}$ satisfies non-stationary ESP under $f|_{\mathcal{S^{\prime}}}$. (b) In the subset ESP, a limited set of vector elements of each state $s\in\mathcal{S}$ has ESP.
@@ II.3 Non-stationary ESP of QRC
((a)) ESP
((b)) Non-stationary ESP
((c)) ESP
((d)) Non-stationary ESP
Figure 4: Grid search result of input rotation axis $\theta$ for a QRC with SK Hamiltonian with external field and qubit-reset for input encoding. Each spherical coordinate corresponds to the input rotation axis $\theta$ described in the main text. (a) ESP characterized by the ESP indicator drawn on the Bloch sphere. (b) Non-stationary ESP characterized by the non-stationary ESP indicator, with values capped at an upper bound of 1, drawn on the Bloch sphere. Our experiments are done with |$|\{\mathbf{u}_{\tau}\}|=200$ and $w=10$, where $w$ is the variable defined in Eq. (2). (c) The flattened surface of the Bloch sphere of traditional ESP in (a). (d) The flattened surface of the Bloch sphere of non-stationary ESP in (d). The top and bottom lines in (c-d) correspond to the north and south poles, respectively. Please note that the strong and weak ESP regions’ pattern is Hamiltonian-dependent. Namely, this result is specific to our selected random Hamiltonian.
To numerically examine the correspondence between non-stationary ESP indicator Eq. (2) and the actual information processing capability of QRC, we prepared a QRC of the following Sherrington–Kirkpatrick (SK) Hamiltonian [25] with an external field:
@@
where $J_{s}=1$, $W=0.312$, and $h=0.013$, in precision up to the third digit after the decimal point.
@@
for each $t$.
@@
((a)) NARMA2 RNMSE
((b)) NARMA10 RNMSE
Figure 5: Root normalized mean squared error (RNMSE) of (a) NARMA2, (b) NARMA10 tasks with the same setup.
@@ II.3.1 ESP and non-stationary ESP
The flattened surface plots in Fig. 4(c) and Fig. 4(d) showed an expanded Bloch sphere surface, where the upper line and bottom line, respectively, correspond to the north and south poles. We can see the difference between ESP and non-stationary ESP in those poles, where ESP is satisfied, while non-stationary ESP is not. Please note that ESP and non-stationary ESP indicators are upper bound by one because indicators greater than one imply that there is no possibility of convergence, even if we increase the input sequence length in the experiments. The boundaries between red and blue regions colored by white are not intentional because we did not cap values by any lower bounds.
These poles correspond to the input configurations in which an input encoding axis direction and the direction of the system state Bloch vector coincide after a reset operation. In such settings, input encoding does not modulate the system’s fixed point, and the system falls into constant output.
@@ II.3.2 NARMA tasks
((a)) X-axis rotation input (Raw)
((b)) X-axis rotation input (LR)
((c)) Y-axis rotation input (Raw)
((d)) Y-axis rotation input (LR)
Figure 6: Raw and linear regression (LR) outputs of 2 different input rotation axis ($\theta$) configurations: (a, b) X-axis and (c, d) Y-axis. (a, c) Raw measurement results. (b, d) Target sequence (blue) and linear regression results (orange).
@@
We observed that the results of both NARMA2 and NARMA10 tasks in Fig. 5(a) and Fig. 5(b), respectively, nicely corresponded to non-stationary ESP results in Fig. 4(d). It should be noted that the red line at the top and bottom of Fig. 4(d), which corresponds to the convergence of the dynamics to a single fixed point, and results in large RNMSE in Fig. 5(a) and Fig. 5(b), cannot be reproduced by the traditional ESP indicator, as shown in Fig. 4(c).
To clarify the actual dynamics and prediction behavior of different input encoding axes in the NARMA2 experiment of Sec. II.3.2, we show actual output signals of two separate input axis configurations depicted in Fig. 6(a) and Fig. 6(c), in which all Pauli string measurement results from $t=3000$ to $t=3100$ are plotted for both setups. In Fig. 6(b) and Fig. 6(d), the target sequence $y$ and the predicted sequence $\hat{y}$ from an optimized linear regression readout for the first $2000$ steps of the test sequence are shown.
In this comparison, the X-axis input demonstrates a configuration where non-stationary ESP does not hold, while the Y-axis input is for non-stationary ESP-compatible demonstration. We cannot identify which has non-stationary ESP by simply seeing the raw output signal in Fig. 6(a) and Fig. 6(c). However, linear readout results in Fig. 6(b) and Fig. 6(d) clearly show the performance difference between the two configurations.
((a)) MC1
((b)) IPCtot
((c)) Linear memory decay ($k\leq 300$)
((d)) Max linear delay
((e)) Rank after washout
((f)) Linear memory decay ($k\leq 10$)
@@ II.3.3 MC and IPC
We observed that the low values of memory function of delay $k=1$ ($MC_{1}$) depicted as the blue region in Fig. 7(a) partially corresponds to the red high-error region in Fig. 5(a). Furthermore, the maximum linear delay, computed by $\mathrm{argmax}_{k}(MC_{k}>10^{-4})$, shown in Fig. 7(d), has global consistency with the non-stationary ESP depicted in Fig. 4(d). However, regions in Fig. 7(d) indicated by black rectangles are inconsistent with the corresponding regions in Fig. 4(d). It can be speculated that those regions have longer memory in non-linear terms, as we can see from the non-saturating $IPC$ values in the region, also indicated by black rectangles in Fig. 7(b).
@@ II.4 Subset non-stationary ESP of QRC
((a)) Total system
((b)) Damping subsystem
((c)) Non-damping subsystem
Figure 9: (a) Non-stationary ESP of the total system, and subset non-stationary ESP of (b) the damping subsystem, and (c) the non-damping subsystem.
@@
Furthermore, we modulated the entanglement between two qubits by applying the matrix power of CNOT gate: $U_{CX}^{p}$ for $p\in[0,1]$ where
@@
Namely, if $p=0$, it equaled an identity, and if $p=1$, it equaled a CNOT gate that took $q_{0}$ as the control qubit. Overall, non-input-driven system dynamics became
@@
((a)) Total system
((b)) Damping subsystem
((c)) Non-damping subsystem
((d)) Total system
((e)) Damping subsystem
((f)) Non-damping subsystem
((g)) Total system
((h)) Damping subsystem
((i)) Non-damping subsystem
((j)) Total system
((k)) Damping subsystem
((l)) Non-damping subsystem
@@ II.4.2 NARMA tasks, MC and IPC
((a)) $MC_{0}$
((b)) $MC_{1}$
((c)) $MC_{2}$
((d)) $MC_{3}$
((e)) $MC_{even}$
((f)) $MC_{odd}$
Figure 11: (a-d) $MC_{0}$ to $MC_{3}$, (e) $MC_{even}$, and (f) $MC_{odd}$ of the damping subsystem.
@@
((a)) Entangling basis
((b)) Total system
Figure 12: $MC_{2+}$ of the entangling basis and the total system.
((a)) $MC_{0}$
((b)) $MC_{1}$
((c)) $MC_{2}$
((d)) $MC_{3}$
((e)) $MC_{4}$
((f)) $MC_{5}$
((g)) $MC_{6}$
((h)) $MC_{7}$
((i)) $MC_{8}$
((j)) $MC_{9}$
((k)) $MC_{10}$
((l)) $MC_{11}$
((m)) $MC_{12}$
((n)) $MC_{13}$
((o)) $MC_{14}$
((p)) $MC_{15}$
((q)) $MC_{16}$
((r)) $MC_{17}$
((s)) $MC_{18}$
((t)) $MC_{19}$
Figure 13: Memory functions of delay $0\leq k\leq 19$ by XZ basis measurement results. The red region has finite memory functions. Please be warned that each heatmap has a different scale. Small values still have statistical meaning because each memory function is thresholded by a random shuffle surrogate [29].

### v5

@@ Abstract
The echo state property (ESP) represents a fundamental concept in the reservoir computing (RC) framework that ensures output-only training of reservoir networks by being agnostic to the initial states and far past inputs. However, the traditional definition of ESP does not describe possible non-stationary systems in which statistical properties evolve. To address this issue, we introduce two new categories of ESP: non-stationary ESP, designed for potentially non-stationary systems, and subspace and subset ESP, designed for systems whose subsystems have ESP. Following the definitions, we numerically demonstrate the correspondence between non-stationary ESP in the quantum reservoir computer (QRC) framework with typical Hamiltonian dynamics and input encoding methods using non-linear autoregressive moving-average (NARMA) tasks. We also confirm the correspondence by computing linear/non-linear memory capacities that quantify input-dependent components within reservoir states. Our study presents a new understanding of the practical design of QRC and other possibly non-stationary RC systems in which non-stationary systems and subsystems are exploited.
@@ I Introduction
In this paper, we define two extensions of traditional ESP. One extension is non-stationary ESP, which requires finite variance output signals relative to initial-state difference decay. Another extension is subset and subspace non-stationary ESP, which focuses on a situation in which a part of the system has a non-stationary ESP, leaving the entire system possibly initial-state dependent. Our analysis includes a numerical study of non-stationary ESP and subset non-stationary ESP in a typical QRC setup with a specific type of Hamiltonian system.
@@
  • Defined non-stationary, subspace, and subset versions of ESP, which could be practical for QRC and other non-conventional systems.
@@ II.1 Non-stationary ESP
((a))
((b))
Figure 2: Schematics of two types of ESPs. State space is illustrated as a circle, and different states are illustrated as red (gray) arrows. (a) Non-stationary ESP. State difference decays, but the state space remains finite. (b) With traditional ESP but not with the non-stationary ESP. It has an input-invariant fixed point. The state difference decays as fast as the state decay. Schematics show representative cases where two states compared in ESP calculation are on the same shrinking state space for simplicity.
@@
$\mathrm{E}_{w}^{t}(\{\mathbf{u}_{\tau}\})$ $\equiv\frac{1}{w}\sum_{k=0}^{w-1}\mathbf{u}_{t-k},$ (3)
$\mathrm{Var}_{w}^{t}(\{\mathbf{u}_{\tau}\})$ $\equiv{\mathrm{E}}_{w}^{t}\left(\{\|\mathbf{u}_{\tau}-\mathrm{E}_{w}^{t}(\{\mathbf{u}_{\tau}\})\|^{2}\}\right),$
$\overline{\mathrm{E}}_{w}^{t}(f;s_{0})$ $\equiv\frac{1}{w}\sum_{k=0}^{w-1}f(\{\mathbf{u}_{\tau}\}_{\tau\leq t-k};s_{0}),$
@@
An example for cases 1 and 2 above would be an input-driven dynamical system with explicit scaling at each time step. For instance, $f$ can be an echo state network [4], and the following two-step state update rule can be written:
[…]
can be converted into the state update rule of $y_{t}$ as below.
$y_{t+1}=c^{t+1}f(\frac{y_{t}}{c^{t}},u_{t}).$ (5)
This will be an example for case 1 if $c\ll 1$ and case 2 if $c\gg 1$, provided that, when $c=1$, it satisfies the non-stationary ESP without shrinking or expanding state space over time. Here, $x_{t}\in\mathbb{R}^{N}$ and $y_{t}\in\mathbb{R}^{N}$ represents reservoir states, and $u_{t}\in\mathbb{R}^{M}$ represents inputs.
An example for case 3 would be an input-driven dynamical system with explicit bias at each time step. For instance, the state update rule will be
[…]
again, provided that when $b=0$, it satisfies the non-stationary ESP without shrinking or expanding state space over time. The equation above can be converted into the single-step state update below.
$y_{t+1}=f(y_{t}-bt,u_{t})+b(t+1).$ (7)
It should be noted that the traditional ESP is also a good indicator of fading memory in this case, although it is not a stationary system.
@@ II.2 Subspace and subset non-stationary ESP
((a))
((b))
Figure 3: (a) Subspace ESP. The full state space $\mathcal{S}$ has an invariant subspace $\mathcal{S^{\prime}}\subseteq\mathcal{S}$ under the input-driven system dynamics $f$. Furthermore, $\mathcal{S^{\prime}}$ satisfies non-stationary ESP under $f|_{\mathcal{S^{\prime}}}$. (b) Subset ESP. A limited set of vector elements of each state $s\in\mathcal{S}$ has ESP.
@@ II.3 Non-stationary ESP of QRC
((a))
((b))
((c))
((d))
Figure 4: Grid search result of input rotation axis $\theta$ for a QRC with SK Hamiltonian with external field and qubit-reset for input encoding. Each spherical coordinate corresponds to the input rotation axis $\theta$ described in the main text. (a) ESP characterized by the ESP indicator drawn on the Bloch sphere. (b) Non-stationary ESP characterized by the non-stationary ESP indicator, drawn on the Bloch sphere. Our experiments are done with |$|\{\mathbf{u}_{\tau}\}|=200$ and $w=10$, where $w$ is the variable defined in Eq. (2). (c) The flattened surface of the Bloch sphere of traditional ESP in (a). (d) The flattened surface of the Bloch sphere of non-stationary ESP in (d). The top and bottom lines in (c-d) correspond to the north and south poles, respectively.
To numerically examine the correspondence between non-stationary ESP and the actual information processing capability of QRC, we prepared a QRC of the following Sherrington–Kirkpatrick (SK) Hamiltonian [25] with an external field:
@@
where $J_{s}=1$, $W=0.312$, and $h=0.013$, in precision up to the third digit after the decimal point. Here $h$ represents constant global external field, and each $D_{i}$ represents random local external field.
@@
for each $t$. Here $I$ is a $2\times 2$ identity matrix, and $X$, $Y$ and $Z$ are Pauli matrices defined as follows:
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad Y=\begin{pmatrix}0&i\\ -i&0\end{pmatrix},\text{ and }Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ (16)
and an example for $P$ would be $I\otimes X\otimes Y\in\mathbb{R}^{8\times 8}$ for a 3-qubit system for instance.
@@
((a))
((b))
Figure 5: Root normalized mean squared error (RNMSE) of (a) NARMA2, (b) NARMA10 tasks. The settings for the Hamiltonian, parametrized input-encoding, and the output signal are the same as the experiment of ESP indicator results in Fig. 4. Please refer to the main text for how we generated input and target signals.
@@ II.3.1 ESP and non-stationary ESP
The flattened surface plots in Fig. 4(c) and Fig. 4(d) showed an expanded Bloch sphere surface, where the upper line and bottom line, respectively, correspond to the north and south poles. We can see the difference between ESP and non-stationary ESP in those poles, where ESP is satisfied, while non-stationary ESP is not. These poles correspond to the input configurations in which an input encoding axis direction and the direction of the system state Bloch vector coincide after a reset operation. In such settings, input encoding does not modulate the system’s fixed point, and the system falls into constant output. That is why the traditional ESP was satisfied at these poles while the non-stationary ESP was not. In simpler words, because qubit reset generates a state that is fixed under input encoding, the system becomes input-independent at those poles.
Please note that ESP and non-stationary ESP indicators greater than one imply that there is no possibility of convergence, even if we increase the input sequence length in the experiments. The boundaries between yellow (light gray) and navy purple (dark gray) regions colored by white are not intentional because we did not cap values by any lower bounds. In addition, the strong and weak ESP regions’ pattern, as shown in Fig. 4, is Hamiltonian-dependent. Namely, this result is specific to our selected random Hamiltonian. However, the following features are invariant under the choice of Hamiltonians:
  1. 1.
    The top and bottom yellow (light gray) lines for the non-stationary ESP, as shown in Fig. 4(d), appear when we use the reset-input encoding because of the reason discussed above.
  2. 2.
    There always exist gradients of strong and weak non-stationary ESP regions. This indicates that the input-encoding method is an important factor in determining the fading memory property of QRC. Please see Appendix B for other examples of the ESP and non-stationary ESP results with different Hamiltonians.
Here, the ESP indicators in Fig. 4(c) are not the same as the largest conditional Lyapunov exponent [26], which determines the diverging or converging behavior of input-driven dynamics around attractors, which can be calculated by numerical methods such as [27]. Because the ESP indicators are not calculated after sufficient washout time, it is not ensured that the input-driven dynamics of the QRC reach an attractor when calculating ESP indicators. Instead, they are calculated for random initial states, which are not guaranteed to be placed near attractors.
However, we argue that the upper bound of the largest Lyapunov exponent in QRC is 0 because any quantum channel is linear and the quantum state space is bounded. It should be noted that this fact also means that there is no classical chaos in quantum systems. It can be expected that the upper bound of zero will be observed in QRC with completely unitary subsystems, including a trivial subsystem of the total system. In our QRC settings described in Sec. 4, completely unitary systems do not appear in any parameter configurations. Therefore, every region in Fig. 4 is expected to have a negative largest Lyapunov exponents.
Because the Lyapunov exponent is always negative in QRC, and we divide the ESP indicator by the variance decay as shown in the definition Eq. (18), there is no simple relationship between the non-stationary ESP indicator and the largest Lyapunov exponent. The non-stationary ESP indicator catches the relative strength of relative divergence between different trajectories and the divergence of each trajectory. Here, we did not calculate the Lyapunov exponents of the QRC systems because of the computational difficulty that arose from the fact that the phase space of the 2-qubit dissipative quantum system is a 16-dimensional complex completely-positive and trace-preserving manifold.
@@ II.3.2 NARMA tasks
((a))
((b))
((c))
((d))
((e))
((f))
Figure 6: Raw (a, b) and linear regression (LR) outputs (c-f) of 2 different input rotation axis ($\theta$) configurations: (a, c, e) X-axis and (b, d, f) Y-axis. (a, b) Raw measurement results. (c, d) Target sequence (dotted blue) and linear regression results (solid orange). (e, f) shows the zoomed signals of (c, d) at time steps $t\in[1000,1100]$ for clarification of the fact that strong non-stationary ESP has positive effects in the NARMA2 task.
@@
We observed that the results of both NARMA2 and NARMA10 tasks in Fig. 5(a) and Fig. 5(b), respectively, nicely corresponded to non-stationary ESP results in Fig. 4(d). It should be noted that the yellow (light gray) line at the top and bottom of Fig. 4(d), which corresponds to the convergence of the dynamics to a single fixed point, and results in large RNMSE in Fig. 5(a) and Fig. 5(b), cannot be reproduced by the traditional ESP indicator, as shown in Fig. 4(c).
To clarify the actual dynamics and prediction behavior of different input encoding axes in the NARMA2 experiment of Sec. II.3.2, we show actual output signals of two separate input axis configurations depicted in Fig. 6(a) and Fig. 6(b), in which all Pauli string measurement results from $t=3000$ to $t=3100$ are plotted for both setups. In Fig. 6(c) and Fig. 6(d), the target sequence $y$ and the predicted sequence $\hat{y}$ from an optimized linear regression readout for the first $2000$ steps of the test sequence are shown.
In this comparison, the X-axis input demonstrates a configuration where non-stationary ESP does not hold, while the Y-axis input is for non-stationary ESP-compatible demonstration. We cannot identify which has non-stationary ESP by simply seeing the raw output signal in Fig. 6(a) and Fig. 6(b). However, linear readout results in Fig. 6(c) and Fig. 6(d) clearly show the performance difference between the two configurations, in the sense that the output signal transformed by the linear readout has better fit to the target signal as shown in Fig. 6(e) and Fig. 6(f).
((a))
((b))
((c))
((d))
((e))
((f))
@@ II.3.3 MC and IPC
We observed that the low values of memory function of delay $k=1$ ($MC_{1}$) depicted as the navy purple (dark gray) region in Fig. 7(a) partially corresponds to the yellow (light gray) high-error region in Fig. 5(a). Furthermore, the maximum linear delay, computed by $\mathrm{argmax}_{k}(MC_{k}>10^{-4})$, shown in Fig. 7(d), has global consistency with the non-stationary ESP depicted in Fig. 4(d). However, regions in Fig. 7(d) indicated by red (gray) rectangles are inconsistent with the corresponding regions in Fig. 4(d). It can be speculated that those regions have longer memory in non-linear terms, as we can see from the non-saturating $IPC$ values in the region, also indicated by red (gray) rectangles in Fig. 7(b).
@@ II.4 Subset non-stationary ESP of QRC
((a))
((b))
((c))
Figure 9: (a) Non-stationary ESP of the total system. Subset non-stationary ESP of (b) the damping subsystem, and (c) the non-damping subsystem.
@@
Furthermore, we modulated the entanglement between two qubits by applying two-qubits unitary that is inspired by a controlled-NOT (CNOT) operation in the QC framework. A controlled-NOT operation is a multiplication of a matrix $U_{CX}$ defined as follows:
@@
For instance, by applying $U_{CX}$, a tensor product density matrix $\rho_{+0}=\begin{pmatrix}\frac{1}{2}&\frac{1}{2}\\ \frac{1}{2}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ becomes
$(U_{CX})\rho_{+0}(U_{CX}^{\dagger})$ (27)
$=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}\begin{pmatrix}\frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&0&0\\ \frac{1}{2}&0&\frac{1}{2}&0\\ 0&0&0&0\\ \end{pmatrix}\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\\ \end{pmatrix}$
$=\begin{pmatrix}\frac{1}{2}&0&0&\frac{1}{2}\\ 0&0&0&0\\ 0&0&0&0\\ \frac{1}{2}&0&0&\frac{1}{2}\\ \end{pmatrix},$
which cannot be written as a tensor product of single qubit density matrices. A two-qubit density matrix that cannot be written as a tensor product of single-qubit density matrices is called entangled. Therefore, the CNOT operation is called an entangling unitary. We modified our entangling unitary by taking a matrix power of $U_{CX}$. We applied $U_{CX}^{p}$ in each time step, where $p$ represents an exponent of the matrix power. Namely, if $p=0$, it equaled an identity, and if $p=1$, it equaled a CNOT operation that took $q_{0}$ as the control qubit. Overall, non-input-driven system dynamics became
@@
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
@@ II.4.2 NARMA tasks, MC and IPC
((a))
((b))
((c))
((d))
((e))
((f))
Figure 11: (a) $MC_{0}$, (b) $MC_{1}$, (c) $MC_{2}$, (d) $MC_{3}$, (e) $MC_{even}$, and (f) $MC_{odd}$ of the damping subsystem.
@@
((a))
((b))
Figure 12: $MC_{2+}$ of (a) the entangling basis, and (b) the total system.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
((m))
((n))
((o))
((p))
((q))
((r))
((s))
((t))
Figure 13: (a-t) Memory functions of delay $0\leq k\leq 19$ by ascending order of $k$, using XZ basis measurement results. The yellow (light gray) region has finite memory functions. Please be warned that each heatmap has a different scale. Small values still have statistical meaning because each memory function is thresholded by a random shuffle surrogate [31].